# Gradient-Mask Tuning Elevates the Upper Limits of LLM Performance

### v1

@@ 4.3 Further Discussions
Figure 3: We employed three distinct strategies for selectively dropping delta parameters at various rates depending on the magnitude (absolute value) of the delta parameters: 1) preferentially dropping salient parameters, 2) preferentially dropping trivial parameters, and 3) dropping parameters randomly.
Our analysis reveals that not all delta parameters contribute equally to the model’s performance. Specifically, we find that parameters with salient values are critical for maintaining the efficacy of the LLM, and their removal leads to a notable and rapid degradation in model functionality. To systematically explore this phenomenon, we designed a series of experiments using the Mistral-7B model, focusing on domain-specific applications in mathematics (GSM8k and MATH). The experimental results are depicted in Figure 3.
Our results indicate substantial differences in the impact of each strategy on model performance. The strategy of dropping a portion of the delta parameters with small absolute values after vanilla fine-tuning demonstrates a relatively robust performance, only showing significant degradation when the dropout rate reaches 80%. This suggests potential benefits from sparsification, which may enhance performance by reducing overfitting or removing noise. In contrast, the random drop strategy leads to an apparent performance decline at a much lower sparsity level of 40%. The strategy of dropping parameters with salient magnitudes results in a precipitous decline in model capabilities. This finding underscores the importance of these significant delta parameters in sustaining the functional integrity of the model. Expectedly, the utilization of the gradient information to trivial drop during the training process can elevate the upper limit of LLM performance and is not constrained by drop rate.
@@
To demonstrate the time-efficiency of proposed GMT, we compare the training speed and computed FLOPs [^6] All data logging is from https://wandb.ai. of GMT, vanilla SFT, and Random Mask, with the metric of number of training samples per second responding to the training speed, and the results are shown in Table 3. The FLOPs computed by the three tuning strategies during training are quite comparable due to the fact that the gradient information utilized by GMT is a by-product of model training and does not impose additional derivation operations. In terms of training speed, the random mask is close to GMT, being 14% and 6% slower than vanilla SFT in the code generation and general domain, respectively. Since GMT needs to calculate the threshold of the gradient being masked based on a preset ratio, a process that requires a TopK operation after the model backpropagates the gradient, and then drops values with smaller absolute values of the gradient based on the threshold. Nonetheless, this trade-off between a small minor time overhead and performance improvement is completely acceptable.
5 Related Works
Delta Tuning (DT) (Hu et al., 2022; Ding et al., 2023) stands out for its efficacy in tuning a subset of the LLM parameters, conserving the integrity of the predominant pre-trained weights. We categorized delta tuning into post-processing and in-training based on the stage of implementation. This section describes the properties of the typical methods of both strategies as well as the respective drawbacks.
Post-processing Strategy
The post-processing strategy, although requiring tuning with full parameter participation, achieves the same purpose as in-training through a series of delta parameter drop and reset strategies. Inspired by dropout, mixout (Lee et al., 2019) stochastically mixes source and target parameters as a regularization technique and extends to fine-tune downstream tasks. DARE (Yu et al., 2023a) finds that setting most (90% or even 99%) of the delta parameters to zero does not diminish the ability to fine-tune the LLM, and extends to model fusion accordingly. Ties-merging (Yadav et al., 2024) describes the delta parameter as a type of task vector that trims it, and then averages the parameters that are identical for multiple model aggregation symbols to achieve model merging. ExPO (Zheng et al., 2024) is inspired by model interpolation (Wortsman et al., 2022) and obtains better aligned models by direct extrapolation from the parameters of the aligned model obtained by DPO or RLHF and the initial SFT model. Although these methods can be implemented directly on fine-tuned models that are open access and require little additional computational overhead, they never achieve optimal performance due to the lack of refinement in the training process.
In-training Strategy
The in-training Strategy is to select a subset of parameters or additional extensions instead of full parameter fine-tuning. The Low-Rank Adaptation (LoRA) (Hu et al., 2021) leverages trainable rank-decomposition matrices to facilitate the fine-tuning process. Prefix-tuning (Li and Liang, 2021) adjusts the PLM by updating a set of prefixed parameters in every transformer layer that are closely related to the task, significantly reducing training costs and enhancing performance across multiple tasks. Adapter (Houlsby et al., 2019; Zhang et al., 2023; Hu et al., 2023) modules add small, trainable networks to each LLM layer, allowing specific fine-tuning while maintaining existing parameters’ stability. Moreover, methods such as Masking (Zhao et al., 2020), DiffPruning (Guo et al., 2021), BitFit (Zaken et al., 2022) and AdaLoRA (Zhang et al., ) have been proposed, which identify a subset of trainable parameters within the LLM to achieve efficient fine-tuning. (Pfeiffer et al., 2023) implements computational units typically as autonomous parameter-efficient modules. Information is routed to a subset of the modules and then aggregated.
These methodologies not only expedite the adaptation process of the model to new tasks but also substantially reduce the consumption of computational resources and training time. In contrast, our approach is focused on the task-specific data information and fine-grained tuning of parameters, which not only preserves the model’s generalizability but also significantly improves its adaptability and performance on specific tasks.
6 Conclusion
@@
In this paper, we propose the Gradient-Mask Tuning (GMT), a pragmatic approach to delta tuning for LLMs. By leveraging gradient magnitudes associated with task-specific data to selectively update parameters, GMT strikes an optimal balance between maintaining computational simplicity and achieving fine-grained parameter tuning. Additionally, GMT identifies more significant parameters and eliminates the redundancy introduced by trivial delta parameters in the fine-tuning. We evaluate the performance of GMT on multiple models in the code generation, math reasoning, and general domain. Experiments demonstrate that GMT not only elevates the upper limits of LLM performance across diverse tasks but also exhibits robustness to various fine-tuning settings. We further analyze the computational efficiency of GMT to confirm that no additional computational FLOPs are needed, as well as acceptable extra time consumption. Moreover, GMT allows for delta tuning without destroying the network structure, and thus can readily replace SFT to accomplish the optimization process.
7 Limitations
This work has several limitations. First, the experiments in this paper could have tried more datasets and more base models. Furthermore, a well-performing LLM requires a sequence of training paradigms, and we have not yet considered the role of GMT in the pre-training and alignment training (e.g., via DPO or RLHF). Finally, this paper has not analyzed the effect of the amount of data on the results, i.e., the comparison of the results of GMT and SFT in the case of limited data.

### v2

@@ Introduction
Our analysis reveals that not all delta parameters contribute equally to the model’s performance. Specifically, we find that parameters with salient values are critical for maintaining the efficacy of the LLM, and their removal leads to a notable and rapid degradation in model functionality. To systematically explore this phenomenon, we designed a series of experiments using the Mistral-7B model, focusing on domain-specific applications in mathematics (GSM8k and MATH). The experimental results are depicted in Appendix.
Our results indicate substantial differences in the impact of each strategy on model performance. The strategy of dropping a portion of the delta parameters with small absolute values after vanilla fine-tuning demonstrates a relatively robust performance, only showing significant degradation when the dropout rate reaches 80%. In contrast, the random drop strategy leads to an apparent performance decline at a much lower sparsity level of 40%. The strategy of dropping parameters with salient magnitudes results in a precipitous decline in model capabilities. This finding underscores the importance of these significant delta parameters in sustaining the functional integrity of the model. Expectedly, the utilization of the gradient information to trivial drop during the training process can elevate the upper limit of LLM performance and is not constrained by drop rate.
@@
To demonstrate the time-efficiency of proposed GMT, we compare the training speed and computed FLOPs of GMT, vanilla SFT, and RMT, with the metric of number of training samples per second responding to the training speed, and the results are shown in Table 3. The FLOPs computed by the three tuning strategies during training are quite comparable due to the fact that the gradient information utilized by GMT is a by-product of model training and does not impose additional derivation operations. In terms of training speed, the RMT is close to GMT, being 14% and 6% slower than vanilla SFT in the code generation and general domain, respectively. Since GMT needs to calculate the threshold of the gradient being masked based on a preset ratio, a process that requires a TopK operation after the model backpropagates the gradient, and then drops values with smaller absolute values of the gradient based on the threshold. Nonetheless, this trade-off between a small minor time overhead and performance improvement is completely acceptable.
Conclusion
@@
In this paper, we propose the Gradient-Mask Tuning (GMT), a pragmatic approach to optimize LLMs by selectively updating parameters based on gradient information. GMT identifies more important parameters to eliminate redundant updates introduced in fine-tuning and improves model performance on various tasks. We evaluate the performance of GMT on multiple models in the code generation, math reasoning, and general domain. Experiments demonstrate that GMT not only elevates the upper limits of LLM performance across diverse tasks but also exhibits robustness to various fine-tuning settings. We further analyze the computational efficiency of GMT to confirm that no additional computational FLOPs are needed, as well as acceptable extra time consumption. Moreover, GMT allows for fine-tuning without destroying the network structure, and thus can readily replace SFT and DPO to accomplish the optimization process.
@@ Appendix B Details of the training mask ratio
Figure 3: We employed three strategies for selectively dropping delta parameters at various rates depending on the magnitude (absolute value) of the delta parameters: 1) preferentially dropping salient parameters, 2) preferentially dropping trivial parameters, and 3) dropping parameters randomly.